# Statistical testing of random number generators and their improvement using randomness extraction

### v1

@@ F.2 Full results
Table 27: Statistical test results for RDSEED as the weak input source to the physical randomness extractor hierarchy level, implemented with the Circulant extractor. The seed is randomness generated using the semi-device-independent randomness amplification protocol outlined in Section F.1. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.
@@
Table 28: Statistical test results for IDQ Quantis as the weak input source to the physical randomness extractor hierarchy level, implemented with the Circulant extractor. The seed is randomness generated using the semi-device-independent randomness amplification protocol outlined in Section F.1. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parenthesis.

### v2

@@ F.2 Full Results
Table 31: Statistical test results for RDSEED as the weak input source to the physical randomness extractor hierarchy level, implemented with the Circulant extractor. The seed is randomness generated using the semi-device-independent randomness amplification protocol outlined in Section F.1. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
Table 32: Cont.
RNG
NIST
Diehard
ENT
SmallCrush
Alphabit
Rabbit
PractRand
@@
Table 33: Statistical test results for IDQ Quantis as the weak input source to the physical randomness extractor hierarchy level, implemented with the Circulant extractor. The seed is randomness generated using the semi-device-independent randomness amplification protocol outlined in Section F.1. In cells with multiple entries, failed tests are on the left and suspicious tests (when applicable) are on the right in parentheses.
@@
{adjustwidth}
[…]